Nevey:

math help plz, screenshots in comments. I did some of the work but im not sure its correct and the very last part i dont understand.

Nevey:

1545140575-5c18f9574da21850752e3400-image.png

Nevey:

1545140615-5c18f9574da21850752e3400-image.png

Nevey:

1545140670-5c18f9574da21850752e3400-image.png

Nevey:

@Kamauri @hero

Kamauri:

that solves the first one

Nevey:

jesus, thats only the 1st one!??!?

Kamauri:

yeah

dude:

Uhhh For the first question, you're supposed to get the same line, not an equation that intersects the line y=2x+1470 is in slope-intercept form but it is not from the equation 2x+3y=1470 How did you get y=2x+1470

Nevey:

i looked up 'slope intercept form' and just kinda went along with the example given .-.

dude:

Well then

Eiwoh2:

^

dude:

When solving to get slope-intercept form you want to isolate x, first you want to do the opposite operation of the "x-term" \(\boxed{2x}+3y=1470\) What is the opposite of +2x?

Nevey:

-2x

dude:

Right so you do that on both sides \(\color{red}{\boxed{2x}-2x}+3y=1470-2x\) The red cancels out because 2-2 is 0 You have \(3y=-2x+1470\) What is the opposite of multiplying?

Nevey:

dividing*

dude:

Right, so we divide by the number in front of the y \(\frac{3y}{\color{red}{3}}=\frac{-2x+1470}{\color{red}{3}}\) The 3's cancel out, so we stay with \(\bf \boxed {y=-\frac{2}{3}+490}\) That is your equation in slope-intercept form

Nevey:

oh, okie. :) ty. ok what about part 2?

dude:

Before we move on, do you know what the slope and y-intercept are here?

Nevey:

im pretty sure the slope is 2/3 and the y intercept is 490 .-. but i cant be 100% sure bc there is no 'y' to give it away

dude:

Good

dude:

So we can use what we know from part 1 to answer part 2

Nevey:

before we move on, this is what i put, correct?

1 attachment
dude:

Yes

Nevey:

ok, now we can move on lmao

Nevey:

we know what the new equation is, as well as the slope and y intercept

dude:

Yes, and I assume you know what the slope and y-intercept do, so could you answer that on your own?

Nevey:

yea i will go ahead and try and when im finished will show u a screenshot

Nevey:

1 attachment
Nevey:

@dude is that correct?

dude:

Good

Nevey:

ok ty